# Recovering Parameters from Edge Fluctuations: Beta-Ensembles and Critically-Spiked Models

### v1

@@ 1.1. Main Result
    Let $\mathbb{F}_{1}=\mathbb{R}$, $\mathbb{F}_{2}=\mathbb{C}$, and $\mathbb{F}_{4}=\mathbb{H}$. If $r>1$ and $\beta\in\{1,2,4\}$, then let $W_{\beta}$ be a standard $r$-dimensional matrix Brownian motion on $\mathbb{F}_{\beta}$ (i.e., with GOE, GUE, or GSE increments depending on whether $\beta=1,2,4$; see [8, Page 2730]), and define
@@
The SAOs were introduced by Edelman-Sutton [24], Ramírez-Rider-Virág [53], and Bloemendal-Virág [7, 8] to describe the edge fluctuations of various point processes of high interest in mathematical physics and statistics. That is, depending on the choice of $\theta$, the eigenvalues of $\mathcal{H}_{\theta}$ either describe the edge fluctuations of a wide class of beta-ensembles with so-called ”regular” external potential [4, 5, 11, 19, 44, 49, 53, 56], or the soft-edge fluctuations of critically-spiked Wishart or Gaussian invariant ensembles with real, complex, or quaternion entries [2, 7, 8, 45, 50, 59].
@@
(1.5) $\mathcal{T}(\Lambda^{\theta})=r_{0}+1/\beta\qquad\text{almost surely.}$.
@@ 1.3. Proof Technique
Our approach to prove Theorem 1.2 is inspired by the theory of inverse problems in spectral geometry. That is, the observation that an impressive amount of information about a differential operator’s domain can be inferred from its eigenvalues alone. One of the most well-known manifestation of this phenomenon is the Weyl law [60] and its various extensions: If $\big{\{}\lambda_{k}(M,b):k\geq 0\big{\}}$ are the eigenvalues of the Laplacian on some compact manifold $M$ subject to some boundary condition $b$ (if $\partial M\neq\varnothing$), then $M$’s dimension, $M$’s volume, $M$’s boundary area, and the boundary condition $b$ (at least Dirichlet versus Neumann) can in many cases be determined by the asymptotics of the eigenvalue counting function
@@
In this context, Question 2.1 can be viewed as the inverse spectral problem for the SAOs, as the components of $\theta=(r,\beta,w)$ characterize $\mathcal{H}_{\theta}$’s domain and noise. Thus, the proof of Theorem 1.2 relies on the observation that $r_{0}+1/\beta$ can be recovered from precise asymptotics of the $\Lambda_{k}^{\theta}$’s as $k\to\infty$. That being said, instead of recovering $r_{0}+1/\beta$ using the asymptotics of the eigenvalue counting function, we do so via the exponential traces $\mathrm{Tr}\big{[}\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big{]}=\sum_{k=0}^{\infty}\mathrm{e}^{-t\Lambda^{\theta}_{k}/2}$ for small $t>0$ (which is also widely used in spectral geometry; see, e.g., [38, Chapter 3] and references therein). More specifically, our proof of Theorem 1.2 relies on showing that
@@
The main technical input in the proof of (1.6) consists of the recently-developed semigroup theory for the multivariate SAOs [29], i.e., the case $r>1$ (which is essential for our application of Theorem 1.2 to the spiked models in Corollary 2.14). This development extended prior results on the semigroup theory of the scalar SAOs [32, 39], i.e., the case $r=1$.
@@ 1.4. Questions of Optimality
More specifically, it is known that for some cases of $\theta$ (notably, $\theta=(1,2,\infty)$; see [31, Proposition 2.27]), the random remainder term $\mathfrak{o}_{t}$ in (1.6) satisfies
@@ 2.1.1. Background
The question of whether the temperature of infinite Gibbs point processes can be recovered from a single sample has attracted some attention in the literature; see, e.g., [20, Section 5.4] and [22] and references therein. In the case of a finite configuration, such as (2.2), this is of course not possible due to the mutual absolute continuity of the particles’ joint densities with different $\beta$’s. However, the problem becomes interesting if one asks about recovering $\beta$ as the number of particles grows to infinity. For instance, several prior works have studied the consistency of maximum likelihood and pseudolikelihood estimators (MLEs and MPLEs) of temperature in spin glasses in the large-dimensional limit; see, e.g., [6, 15, 17, 33] and references therein. In view of extending this to beta-ensembles, one may then ask:
@@
To the best of our knowledge, this has not been studied in the literature. That being said, one of the main hurdles in applying MLEs to Gibbs point processes consists of computing the partition function (e.g., [61, Section 1]). Therefore, given that $Z_{\beta,n}$ is very well understood for beta-ensembles—see [9, 10] and references therein—we expect that a similar strategy could be used to solve Question 2.1. To make this more concrete, consider the following heuristic, (which is informal as the asymptotics therein were calculated using Wolfram Mathematica; see Appendix A):
@@
As argued in Appendix A, the limit (2.4) relies on an asymptotic analysis of the partition function. Thus, we expect that (2.4) could be extended to more general $V$’s using the asymptotic expansions of general beta-ensemble partition functions in [9, (1-3)] and [10, (1.14)].
@@ 2.1.2. New Result
A common feature of the statistics discussed above (i.e., MLEs, MPLEs, or the total energy in (2.3)) is that they use every point in the configuration $\mathcal{B}^{\beta,n}_{1}\leq\cdots\leq\mathcal{B}^{\beta,n}_{n}$ as an input. In sharp contrast to this, Theorem 1.2 implies that the temperature of many beta-ensembles can be recovered from a much more parsimonious asymptotic, which only depends on the vanishing proportion of particles that contribute to edge scaling limits (as the vast majority of particles instead contribute to the bulk scaling limits; e.g., [57, 58]). This is based on the edge universality of beta-ensembles, part of which can be informally summarized as follows:
@@
Let $V$ be a regular one-cut potential. That is, $V$ has an equilibrium measure $\mu_{V}$ supported on a compact interval $[\mathfrak{a},\mathfrak{b}]$, and $\mu_{V}$’s density vanishes like a square root at $\mathfrak{a}$ and $\mathfrak{b}$. Under additional technical conditions on $V$, for any edge $\mathfrak{e}\in\{\mathfrak{a},\mathfrak{b}\}$, there exists scaling constants $\mathfrak{s}_{n}=O(n^{2/3})$ such that
@@
See [4, 11, 19, 44, 49, 53, 56] for the details, as well as [5] for a multi-cut extension of Theorem 2.3, whose statement we omit for simplicity. Given that $r_{0}=0$ when $\theta=(1,\beta,\infty)$, we thus obtain the following corollary from Theorem 1.2:
@@ 2.1.3. Open Problem
for $\varepsilon_{n},t_{n}=o(1)$. A similar result already exists in the special case where $V(x)=x^{2}/2$ when $t_{n}$ is replaced by a fixed $t>0$; see the proof of [39, Corollary 2.10].
@@ 2.2.1. Background
In the early 2000’s, Johnstone introduced the spiked Wishart ensemble [42, Pages 301–304] as a means of modeling the effect of low-rank signals on the spectral statistics of covariance estimators in high dimension. Some years later, Péché [50] introduced a closely related additive spiked ensemble to study low-rank perturbations of self-adjoint matrices. In their simplest incarnation (i.e., rotationally-symmetric noise), these models are defined as follows: Let $r\in\mathbb{N}$, $\beta=\{1,2,4\}$, and $\ell=(\ell_{r-1},\ldots,\ell_{0})\in[0,\infty)^{r}$ be fixed, assuming that $\ell_{r-1}\leq\cdots\leq\ell_{0}$.
@@
(2.7) $p(n)=\gamma\big{(}1+o(1)\big{)}\qquad\text{as }n\to\infty$
@@
The first papers in this direction were by Baik-Ben Arous-Péché [2] and Péché [50]. Among other things, these results showed that the almost-sure asymptotic detectability of each signal $\ell_{i}$ depends on its size relative to the threshold
@@
    The supercritical case, where $\ell_{i}>\tau$ for at least one $10\leq i\leq r-$ 1;
@@
Theorem 2.7 ([2, 3, 14, 50]).
@@
(2.9) $\lim_{n\to\infty}\mathcal{S}^{\ell,\beta,n}_{n-i}=\begin{cases}\mathcal{O}(\ell_{i})&\text{if $i\leq r-1$ and $\ell_{i}>\tau$},\\ \mathfrak{e}&\text{if $i\leq r-1$ and $\ell_{i}\leq\tau$, or if $i>r-1$}\end{cases}\qquad\text{almost surely}.$
@@
Theorem 2.9 ([41, 46, 47, 48]).
@@
(2.10) $\boldsymbol{H}_{0}\text{ : "no perturbation"}\qquad\text{versus}\qquad\boldsymbol{H}_{\ell}\text{ : "perturbation $=\ell$"}$ ”
with nontrivial statistical power when $\ell\neq 0$ is subcritical. The short answer to this question is yes; see, e.g., [25, 41, 43, 47, 48] and references therein for a non-exhaustive sample of this very sizeable literature. A unifying feature of all these results is that the statistical tests that achieve optimal or close-to-optimal power in distinguishing (2.10) (e.g., likelihood ratios or linear spectral statistics) all use the full spectrum $\mathcal{S}^{\beta,\ell,n}_{1}\leq\cdots\leq\mathcal{S}^{\beta,\ell,n}_{n}$ as an input. In the words of Dobriban [23]:
@@
(2.12) $\lim_{n\to\infty}\mathcal{E}_{n}(\ell)=\mathrm{erfc}\left(\tfrac{r}{4}\sqrt{-\log(1-\lambda)}\right)\in(0,1),$
@@
Thirdly, we did not find any works in the literature that directly address Question 2.6 in the critical case $\ell_{r_{\mathrm{c}}-1},\ldots,\ell_{0}=\tau+o(1)$. However, as pointed out in [51, Appendix A], consistent estimators of $\ell$ in this case can be proved to exist by examining the asymptotic power of subcritical tests when $\ell_{i}\to\tau^{-}$ (e.g., (2.12) goes to zero when $\ell_{i}=\lambda\to 1^{-}$). Thus, the answer to Question 2.6 is positive in the critical case. However, to the best of our knowledge, since this conclusion relies on the optimal subcritical estimators, it follows from Remark 2.10 that all eigenvalues are used in the currently-known methods to consistently recover critical signals.
@@ 2.2.2. New Result
Theorem 2.13 ([7, 8]; see also [2, 45, 50, 59]).
@@
exists and is an element of (,]($\infty$,$\infty$]( - ∞ , ∞ ]; then let $w=(w_{1},\ldots,w_{r})$. If we define
@@ 2.2.3. Open Problem
for fixed $t>0$ has not been proved for any spiked model when $r_{0}>0$, even in the simplest case of the scalar SAO (i.e., $r=r_{0}=1$). Indeed, using current techniques, such a result would first require proving a seemingly difficult strong invariance principle for reflected random walks; see [27, Conjectures 2.23 and 6.11].
@@ 2.3.1. Background
In the last 15 years, a sizeable literature on ”rigidity” properties of point processes has been developing, starting with pioneering works by Ghosh and Peres [34, 37]. We refer to [18, 35] and references therein for a survey. In this literature, rigidity refers to the observation that for certain strongly-correlated point processes, some features of the point configuration inside a bounded set is almost-surely determined by the configuration of points outside that set. For instance:
@@
The Ginibre point process was shown to be number rigid in [37].
@@
The $\alpha$-Gaussian analytic function ($\alpha$-GAF) zero point process (i.e., the zeros of the random entire function $z\mapsto\sum_{k\geq 0}\frac{g_{k}}{(k!)^{\alpha/2}}z^{k}$, where $g_{k}$ are i.i.d. complex Gaussians) were shown to be $k$-moment rigid for all $0\leq k\leq\lfloor 1/\alpha\rfloor$ in [36, 37].
@@
Rigidity was in part introduced to help distinguish structured point processes that are otherwise difficult to tell apart. As stated in [37, Page 1794]:
@@
    For the Ginibre process, it was shown in [37] that the number of points is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
@@
    For the $\alpha$-GAF zeros, it was shown in [26, 37] that the $k^{\mathrm{th}}$ moments for $0\leq k\leq\lfloor 1/\alpha\rfloor$ is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
A formal version of these optimality results can be stated using the notion of tolerance; see, e.g., [36, Definitions 2 and 3].
Going back to our main topic, the number rigidity of $\Lambda^{\theta}$ was proved in a sequence of three papers. Namely, first for $\theta=(1,2,\infty)$ in [12], then for $\theta=(1,\beta,w)$ with any choice of $\beta>0$ and $w\in(-\infty,\infty]$ in [30], and finally for all $\theta\in\Theta$ in [29].

### v2

@@ 1.1. Main Result
    Let $\mathbb{F}_{1}=\mathbb{R}$, $\mathbb{F}_{2}=\mathbb{C}$, and $\mathbb{F}_{4}=\mathbb{H}$. If $r>1$ and $\beta\in\{1,2,4\}$, then let $W_{\beta}$ be a standard $r$-dimensional matrix Brownian motion on $\mathbb{F}_{\beta}$ (i.e., with GOE, GUE, or GSE increments depending on whether $\beta=1,2,4$; see [BloemendalVirag2, Page 2730]), and define
@@
The SAOs were introduced by Edelman-Sutton [EdelmanSutton], Ramírez-Rider-Virág [RamirezRiderVirag], and Bloemendal-Virág [BloemendalVirag1, BloemendalVirag2] to describe the edge fluctuations of various point processes of high interest in mathematical physics and statistics. That is, depending on the choice of $\theta$, the eigenvalues of $\mathcal{H}_{\theta}$ either describe the edge fluctuations of a wide class of beta-ensembles with so-called ”regular” external potential [BekermanFigalliGuionnet, Bekerman, BourgadeErdosYau, DeiftGioev, KrishnapurRiderVirag, PasturShcherbina, RamirezRiderVirag, Shcherbina], or the soft-edge fluctuations of critically-spiked Wishart or Gaussian invariant ensembles with real, complex, or quaternion entries [BaikBenArousPeche, BloemendalVirag1, BloemendalVirag2, Mo, Peche, Wang].
@@
(1.5) $\mathcal{T}(\Lambda^{\theta})=r_{0}+1/\beta\qquad\text{almost surely.}$
@@ 1.3. Proof Technique
Our approach to prove Theorem 1.2 is inspired by the theory of inverse problems in spectral geometry. That is, the observation that an impressive amount of information about a differential operator’s domain can be inferred from its eigenvalues alone. One of the most well-known manifestation of this phenomenon is the Weyl law [Weyl] and its various extensions: If $\big\{\lambda_{k}(M,b):k\geq 0\big\}$ are the eigenvalues of the Laplacian on some compact manifold $M$ subject to some boundary condition $b$ (if $\partial M\neq\varnothing$), then $M$’s dimension, $M$’s volume, $M$’s boundary area, and the boundary condition $b$ (at least Dirichlet versus Neumann) can in many cases be determined by the asymptotics of the eigenvalue counting function
@@
In this context, Question 2.1 can be viewed as the inverse spectral problem for the SAOs, as the components of $\theta=(r,\beta,w)$ characterize $\mathcal{H}_{\theta}$’s domain and noise. Thus, the proof of Theorem 1.2 relies on the observation that $r_{0}+1/\beta$ can be recovered from precise asymptotics of the $\Lambda_{k}^{\theta}$’s as $k\to\infty$. That being said, instead of recovering $r_{0}+1/\beta$ using the asymptotics of the eigenvalue counting function, we do so via the exponential traces $\mathrm{Tr}\big[\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big]=\sum_{k=0}^{\infty}\mathrm{e}^{-t\Lambda^{\theta}_{k}/2}$ for small $t>0$ (which is also widely used in spectral geometry; see, e.g., [Gilkey, Chapter 3] and references therein). More specifically, our proof of Theorem 1.2 relies on showing that
@@
The main technical input in the proof of (1.6) consists of the recently-developed semigroup theory for the multivariate SAOs [RigidityMSAO], i.e., the case $r>1$ (which is essential for our application of Theorem 1.2 to the spiked models in Corollary 2.14). This development extended prior results on the semigroup theory of the scalar SAOs [GaudreauLamarreShkolnikov, GorinShkolnikov], i.e., the case $r=1$.
@@ 1.4. Questions of Optimality
More specifically, it is known that for some cases of $\theta$ (notably, $\theta=(1,2,\infty)$; see [RigidityFK, Proposition 2.27]), the random remainder term $\mathfrak{o}_{t}$ in (1.6) satisfies
@@ 2.1.1. Background
The question of whether the temperature of infinite Gibbs point processes can be recovered from a single sample has attracted some attention in the literature; see, e.g., [Dereudre, Section 5.4] and [DereudreLavancier] and references therein. In the case of a finite configuration, such as (2.2), this is of course not possible due to the mutual absolute continuity of the particles’ joint densities with different $\beta$’s. However, the problem becomes interesting if one asks about recovering $\beta$ as the number of particles grows to infinity. For instance, several prior works have studied the consistency of maximum likelihood and pseudolikelihood estimators (MLEs and MPLEs) of temperature in spin glasses in the large-dimensional limit; see, e.g., [BhattacharyaMukherjee, Chatterjee, CometsGidas, GhosalMukherjee] and references therein. In view of extending this to beta-ensembles, one may then ask:
@@
To the best of our knowledge, this has not been studied in the literature. That being said, one of the main hurdles in applying MLEs to Gibbs point processes consists of computing the partition function (e.g., [Younes, Section 1]). Therefore, given that $Z_{\beta,n}$ is very well understood for beta-ensembles—see [BorotGuionnet1, BorotGuionnet2] and references therein—we expect that a similar strategy could be used to solve Question 2.1. To make this more concrete, consider the following heuristic (which is informal as the asymptotics therein were calculated using Wolfram Mathematica; see Appendix A):
@@
As argued in Appendix A, the limit (2.4) relies on an asymptotic analysis of the partition function. Thus, we expect that (2.4) could be extended to more general $V$’s using the asymptotic expansions of general beta-ensemble partition functions in [BorotGuionnet1, (1-3)] and [BorotGuionnet2, (1.14)].
@@ 2.1.2. New Result
A common feature of the statistics discussed above (i.e., MLEs, MPLEs, or the total energy in (2.3)) is that they use every point in the configuration $\mathcal{B}^{\beta,n}_{1}\leq\cdots\leq\mathcal{B}^{\beta,n}_{n}$ as an input. In sharp contrast to this, Theorem 1.2 implies that the temperature of many beta-ensembles can be recovered from a much more parsimonious asymptotic, which only depends on the vanishing proportion of particles that contribute to edge scaling limits (as the vast majority of particles instead contribute to the bulk scaling limits; e.g., [ValkoVirag, ValkoVirag2]). This is based on the edge universality of beta-ensembles, part of which can be informally summarized as follows:
@@
Let $V$ be a regular one-cut potential. That is, $V$ has an equilibrium measure $\mu_{V}$ supported on a compact interval $\mathfrak{a},\mathfrak{b}]$, and $\mu_{V}$’s density vanishes like a square root at $\mathfrak{a}$ and $\mathfrak{b}$. Under additional technical conditions on $V$, for any edge $\mathfrak{e}\in\{\mathfrak{a},\mathfrak{b}\}$, there exists scaling constants $\mathfrak{s}_{n}=O(n^{2/3})$ such that
@@
See [BekermanFigalliGuionnet, BourgadeErdosYau, DeiftGioev, KrishnapurRiderVirag, PasturShcherbina, RamirezRiderVirag, Shcherbina] for the details, as well as [Bekerman] for a multi-cut extension of Theorem 2.3, whose statement we omit for simplicity. Given that $r_{0}=0$ when $\theta=(1,\beta,\infty)$, we thus obtain the following corollary from Theorem 1.2:
@@ 2.1.3. Open Problem
for $\varepsilon_{n},t_{n}=o(1)$. A similar result already exists in the special case where $V(x)=x^{2}/2$ when $t_{n}$ is replaced by a fixed $t>0$; see the proof of [GorinShkolnikov, Corollary 2.10].
@@ 2.2.1. Background
In the early 2000’s, Johnstone introduced the spiked Wishart ensemble [Johnstone, Pages 301–304] as a means of modeling the effect of low-rank signals on the spectral statistics of covariance estimators in high dimension. Some years later, Péché [Peche] introduced a closely related additive spiked ensemble to study low-rank perturbations of self-adjoint matrices. In their simplest incarnation (i.e., rotationally-symmetric noise), these models are defined as follows: Let $r\in\mathbb{N}$, $\beta\in\{1,2,4\}$, and $\ell=(\ell_{r-1},\ldots,\ell_{0})\in[0,\infty)^{r}$ be fixed, assuming that $\ell_{r-1}\leq\cdots\leq\ell_{0}$.
@@
(2.7) $p(n)=\gamma n\big(1+o(1)\big)\qquad\text{as }n\to\infty$
@@
The first papers in this direction were by Baik-Ben Arous-Péché [BaikBenArousPeche] and Péché [Peche]. Among other things, these results showed that the almost-sure asymptotic detectability of each signal $\ell_{i}$ depends on its size relative to the threshold
@@
    The supercritical case, where $\ell_{i}>\tau$ for at least one $0\leq i\leq r-1$;
@@
Theorem 2.7 ([BaikBenArousPeche, BaikSilverstein, CapitaineDonatiMartinFeral, Peche]).
@@
[…]
@@
Theorem 2.9 ([JungChungLee, MontanariReichmanZeitouni, OnatskiMoreiraHallin, OnatskiMoreiraHallin2]).
@@
[…]
with nontrivial statistical power when $\ell\neq 0$ is subcritical. The short answer to this question is yes; see, e.g., [ElAlaouiKrzakalaJordan, JungChungLee, JohnstoneOnatski, OnatskiMoreiraHallin, OnatskiMoreiraHallin2] and references therein for a non-exhaustive sample of this very sizeable literature. A unifying feature of all these results is that the statistical tests that achieve optimal or close-to-optimal power in distinguishing (2.10) (e.g., likelihood ratios or linear spectral statistics) all use the full spectrum $\mathcal{S}^{\beta,\ell,n}_{1}\leq\cdots\leq\mathcal{S}^{\beta,\ell,n}_{n}$ as an input. In the words of Dobriban [Dobriban]:
@@
(2.12) $\lim_{n\to\infty}\mathcal{E}_{n}(\lambda)=\mathrm{erfc}\left(\tfrac{r}{4}\sqrt{-\log(1-\lambda)}\right)\in(0,1),$
@@
Thirdly, we did not find any works in the literature that directly address Question 2.6 in the critical case $\ell_{r_{\mathrm{c}}-1},\ldots,\ell_{0}=\tau+o(1)$. However, as pointed out in [PerryWeinBandeiraMoitra, Appendix A], consistent estimators of $\ell$ in this case can be proved to exist by examining the asymptotic power of subcritical tests when $\ell_{i}\to\tau^{-}$ (e.g., (2.12) goes to zero when $\ell_{i}=\lambda\to 1^{-}$). Thus, the answer to Question 2.6 is positive in the critical case. However, to the best of our knowledge, since this conclusion relies on the optimal subcritical estimators, it follows from Remark 2.10 that all eigenvalues are used in the currently-known methods to consistently recover critical signals.
@@ 2.2.2. New Result
Theorem 2.13 ([BloemendalVirag1, BloemendalVirag2]; see also [BaikBenArousPeche, Mo, Peche, Wang]).
@@
exists and is an element of $(-\infty,\infty]$; then let $w=(w_{1},\ldots,w_{r})$. If we define
@@ 2.2.3. Open Problem
for fixed $t>0$ has not been proved for any spiked model when $r_{0}>0$, even in the simplest case of the scalar SAO (i.e., $r=r_{0}=1$). Indeed, using current techniques, such a result would first require proving a seemingly difficult strong invariance principle for reflected random walks; see [ConvergenceToSpiked, Conjectures 2.23 and 6.11].
@@ 2.3.1. Background
In the last 15 years, a sizeable literature on ”rigidity” properties of point processes has been developing, starting with pioneering works by Ghosh and Peres [Ghosh, GhoshPeres]. We refer to [Coste, GhoshLebowitz] and references therein for a survey. In this literature, rigidity refers to the observation that for certain strongly-correlated point processes, some features of the point configuration inside a bounded set is almost-surely determined by the configuration of points outside that set. For instance:
@@
The Ginibre point process was shown to be number rigid in [GhoshPeres].
@@
The $\alpha$-Gaussian analytic function ($\alpha$-GAF) zero point process (i.e., the zeros of the random entire function $z\mapsto\sum_{k\geq 0}\frac{g_{k}}{(k!)^{\alpha/2}}z^{k}$, where $g_{k}$ are i.i.d. complex Gaussians) were shown to be $k$-moment rigid for all $0\leq k\leq\lfloor 1/\alpha\rfloor$ in [GhoshKrishnapur, GhoshPeres].
@@
Rigidity was in part introduced to help distinguish structured point processes that are otherwise difficult to tell apart. As stated in [GhoshPeres, Page 1794]:
@@
    For the Ginibre process, it was shown in [GhoshPeres] that the number of points is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
@@
    For the $\alpha$-GAF zeros, it was shown in [GangopadhyayGhoshTan, GhoshPeres] that the $k^{\mathrm{th}}$ moments for $0\leq k\leq\lfloor 1/\alpha\rfloor$ is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
A formal version of these optimality results can be stated using the notion of tolerance; see, e.g., [GhoshKrishnapur, Definitions 2 and 3].
Going back to our main topic, the number rigidity of $\Lambda^{\theta}$ was proved in a sequence of three papers. Namely, first for $\theta=(1,2,\infty)$ in [Bufetov16], then for $\theta=(1,\beta,w)$ with any choice of $\beta>0$ and $w\in(-\infty,\infty]$ in [RigiditySAO], and finally for all $\theta\in\Theta$ in [RigidityMSAO].